# 3D Diffusion Policy: Generalizable Visuomotor Policy Learning via Simple 3D Representations

### v3

@@ Abstract
Imitation learning provides an efficient way to teach robots dexterous skills; however, learning complex skills robustly and generalizablely usually consumes large amounts of human demonstrations. To tackle this challenging problem, we present 3D Diffusion Policy (DP3), a novel visual imitation learning approach that incorporates the power of 3D visual representations into diffusion policies, a class of conditional action generative models. The core design of DP3 is the utilization of a compact 3D visual representation, extracted from sparse point clouds with an efficient point encoder. In our experiments involving 72 simulation tasks, DP3 successfully handles most tasks with just 10 demonstrations and surpasses baselines with a 24.2% relative improvement. In 4 real robot tasks, DP3 demonstrates precise control with a high success rate of 85%, given only 40 demonstrations of each task, and shows excellent generalization abilities in diverse aspects, including space, viewpoint, appearance, and instance. Interestingly, in real robot experiments, DP3 rarely violates safety requirements, in contrast to baseline methods which frequently do, necessitating human intervention. Our extensive evaluation highlights the critical importance of 3D representations in real-world robot learning. Videos, code, and data are available on 3d-diffusion-policy.github.io.
Figure 1: 3D Diffusion Policy (DP3) is a universal visual imitation learning algorithm that marries 3D visual representations with diffusion policies, achieving surprising effectiveness in diverse simulated and real-world tasks, including both high-dimensional and low-dimensional control tasks, with a practical inference speed.
@@ I Introduction
Imitation learning provides an efficient way to teach robots a wide range of motor skills, such as grasping [69, 61, 83], legged locomotion [42], dexterous manipulation [1, 17, 11], humanoid loco-manipulation [55], and mobile manipulation [58, 13]. Visual imitation learning, which takes high-dimensional visual observations such as images or depth maps, eases the need for task-specific state estimation and thus gains the popularity [10, 61, 83, 12, 21].
@@
    We conduct comprehensive analyses on visual representations in DP3 and show that a simple point cloud representation might be preferred over other intricate 3D representations.
@@
    Our real-world experiments demonstrate that complex high-dimensional dexterous manipulation tasks could be handled with simple design and little human data.
DP3 emphasizes the power of marrying 3D representations with diffusion policies in real-world robot learning. Videos, code, and data are available on 3d-diffusion-policy.github.io.
@@ II-A Diffusion Models in Robotics
Diffusion models, a category of generative models that progressively transform random noise into a data sample, have achieved great success in high-fidelity image generation [24, 64, 52, 63]. Owing to their impressive expressiveness, diffusion models have recently been applied in robotics, including in fields such as reinforcement learning [71, 2], imitation learning [10, 41, 51, 73, 65], reward learning [26, 39], grasping [72, 67, 62], and motion planning [53, 28]. In this work, we focus on representing visuomotor policies as conditional diffusion models, referred to as diffusion policies, following the framework established in [10, 41]. Unlike prior methods that primarily focus on images and states as conditions, we pioneer in incorporating 3D conditioning into diffusion policies.
@@ II-B Visual Imitation Learning
Figure 2: Overview of 3D Diffusion Policy (DP3). Above: In the training phase, DP3 simultaneously trains its perception module and decision-making module in an end-to-end manner using expert demonstrations. During evaluation, DP3 determines actions based on visual observations from the environment. Below: DP3 perceives its environment through single-view point clouds. The sparsely sampled point clouds are encoded into compact 3D representations by a lightweight MLP encoder. Subsequently, DP3 generates actions conditioning on these 3D representations and the robot states, using a diffusion model as the backbone.
@@ II-C Learning Dexterous Skills
Achieving human-like manipulation skills in robots has been a longstanding objective pursued by robotics researchers. Reinforcement learning has been a key tool in this endeavor, enabling robots with dexterous hands to master a variety of tasks, such as pouring water [48, 82], opening doors [50, 22, 8], rotating objects [45, 77, 79, 46], reorienting objects [19, 7, 6], spinning pens [35], grasping tools [1], executing handovers [84, 25], and building Legos [9]. Imitation learning offers another pathway, with approaches like DIME [3], DexMV [48], and DexCap [70] translating human hand movements into robotic actions through retargeting and enabling learning from humans. Our work, however, diverges from these specific design-centric methods. We demonstrate that enabling the acquisition of these complex skills with minimal demonstrations could be achieved by improving the imitation learning algorithm itself.
@@ III-C Decision
Implementation details. We use the convolutional network-based diffusion policy [10]. We use DDIM [63] as the noise scheduler and use sample prediction instead of epsilon prediction for better high-dimensional action generation, with 100 timesteps at training and 10 timesteps at inference. We train 1000 epochs for MetaWorld tasks and 3000 epochs for other simulated and real-world tasks, with batch size 128 for DP3 and all the baselines.
@@ IV-B Efficiency and Effectiveness
    High accuracy. Summarized results are in Figure 1(a) and results for each domain are in Table I. We observe that DP3 achieves a success rate 100% in nearly 20 tasks, whereas Diffusion Policy does in around 10 tasks. Additionally, we observe that Diffusion Policy records more than 20 tasks with a success rate lower than 30% given a small set of demonstrations, while DP3 could generally achieve higher accuracy.
@@
    Learning efficiency. While we train enough epochs to guarantee convergence, we observe that DP3 typically reaches convergence within approximately 500 epochs across all tasks, as illustrated in Figure 5. In contrast, Diffusion Policy tends to converge at a much slower pace or converge into sub-optimal results.
@@
    Efficient scaling with demonstrations. As shown in Figure 6, we find that in Adroit tasks, both DP3 and Diffusion Policy perform reasonably while DP3 achieves a comparable accuracy with fewer demonstrations. For MetaWorld tasks above the easy level, Diffusion Policy fails to learn even with an increased number of demonstrations, lagging significantly behind DP3. This underscores that the 3D modality is not just beneficial but essential for certain manipulation tasks.
@@ IV-C Ablations
Choice of 3D representations. In DP3, we deliberately select point clouds to represent the 3D scene. To compare different choices of 3D representations, we implement other 3D representations, including RGB-D, depth, and voxel. The RGB-D and depth images are processed using the same image encoder as Diffusion Policy, while voxel representations employ the VoxelCNN, as implemented in [7]. As demonstrated in Table IV, these alternative 3D representations fall short of DP3. Notably, RGB-D and depth images are very close and not comparable to point clouds, indicating that the proper usage of depth information is essential.
@@
Point cloud $\mathbf{100\scriptstyle{\pm 0}}$ $\mathbf{62\scriptstyle{\pm 4}}$ $\mathbf{43\scriptstyle{\pm 6}}$ $\mathbf{99\scriptstyle{\pm 1}}$ $\mathbf{69\scriptstyle{\pm 4}}$ $\mathbf{97\scriptstyle{\pm 4}}$ $\mathbf{78.3}$
@@
Design choices in DP3. Besides the 3D representations, the effectiveness of DP3 is contributed by several small design choices, as shown in Table VII. (a) Cropping point clouds helps largely improve accuracy; (b) Incorporating LayerNorm layers could help stabilize training across different tasks [23, 4]; (c) Sample prediction in the noise sampler brings faster convergence, also shown in Figure 7; (d) The projection head in DP3 Encoder accelerates the inference by projecting features to the lower dimension, without hurting accuracy; (e) Removing color channels ensures robust appearance generalization; (f) In low-dimensional control tasks, DPM-solver++ [34] as the noise sampler is competitive to DDIM, while DPM-solver++ could not handle high-dimensional control tasks well empirically.
@@ V-C Generalization
View generalization. Generalizing image-based methods across different views is notably challenging [76], and acquiring training data from multiple views can be time-consuming and costly [83, 59]. We demonstrate in Table XII that DP3 effectively addresses this generalization problem when the camera views are altered slightly. It is important to note that since the camera view is altered, we manually adjust the cropped space of the point clouds.
@@
V-D An Interesting Observation: Safe Deployment
In our real-world experiments, we observe that image-based and depth-based diffusion policies often deliver unpredictable behaviors in real-world experiments, which necessitates human termination to ensure robot safety. We define this situation as safety violation and compute the safety violation rate in our main real-world experiments, shown in Table XIII. Interestingly and surprisingly, we find that DP3 rarely violates the safety, showing that DP3 is a practical and hardware-friendly method for real robot learning. An intuitive explanation is that since the robots operate in 3D space, directly observing 3D information helps avoid collision. It is important to note that our assessment of safety is primarily qualitative. We intend to explore a more theoretical understanding of this observation in our future work.
TABLE XIII: Safety violation rate. While conducting the main real-world experiments, we also count the times of safety violation and compute the rate.
@@ Acknowledgement
We would like to thank Zhecheng Yuan, Cheng Lu, and Jianfei Chen for their helpful discussions. This work is supported by National Key R&D Program of China (2022ZD0161700).

### v4

@@ Abstract
Imitation learning provides an efficient way to teach robots dexterous skills; however, learning complex skills robustly and generalizablely usually consumes large amounts of human demonstrations. To tackle this challenging problem, we present 3D Diffusion Policy (DP3), a novel visual imitation learning approach that incorporates the power of 3D visual representations into diffusion policies, a class of conditional action generative models. The core design of DP3 is the utilization of a compact 3D visual representation, extracted from sparse point clouds with an efficient point encoder. In our experiments involving 72 simulation tasks, DP3 successfully handles most tasks with just 10 demonstrations and surpasses baselines with a 24.2% relative improvement. In 4 real robot tasks, DP3 demonstrates precise control with a high success rate of 85%, given only 40 demonstrations of each task, and shows excellent generalization abilities in diverse aspects, including space, viewpoint, appearance, and instance. Interestingly, in real robot experiments, DP3 rarely violates safety requirements, in contrast to baseline methods which frequently do, necessitating human intervention. Our extensive evaluation highlights the critical importance of 3D representations in real-world robot learning. Code and videos are available on 3d-diffusion-policy.github.io .
Figure 1: 3D Diffusion Policy (DP3) is a visual imitation learning algorithm that marries 3D visual representations with diffusion policies, achieving surprising effectiveness in diverse simulation and real-world tasks, with a practical inference speed.
@@ I Introduction
Imitation learning provides an efficient way to teach robots a wide range of motor skills, such as grasping [69, 61, 83], legged locomotion [41], dexterous manipulation [1, 16], humanoid loco-manipulation [55], and mobile manipulation [58, 12]. Visual imitation learning, which takes high-dimensional visual observations such as images or depth maps, eases the need for task-specific state estimation and thus gains the popularity [10, 61, 83, 11, 20].
@@
    We conduct comprehensive analyses on visual representations in DP3 and show that a simple point cloud representation is preferred over other intricate 3D representations and is better suited for diffusion policies over other policy backbones.
@@
    DP3 is able to perform real-world deformable object manipulation using a dexterous hand with only 40 demonstrations, demonstrating that complex high-dimensional tasks could be handled with little human data.
DP3 emphasizes the power of marrying 3D representations with diffusion policies in real-world robot learning. Code is available on https://github.com/YanjieZe/3D-Diffusion-Policy.
@@ II-A Diffusion Models in Robotics
Diffusion models, a category of generative models that progressively transform random noise into a data sample, have achieved great success in high-fidelity image generation [23, 64, 52, 63]. Owing to their impressive expressiveness, diffusion models have recently been applied in robotics, including in fields such as reinforcement learning [71, 2], imitation learning [10, 40, 51, 73, 65, 42], reward learning [25, 38], grasping [72, 67, 62], and motion planning [53, 27]. In this work, we focus on representing visuomotor policies as conditional diffusion models, referred to as diffusion policies, following the framework established in [10, 40]. Unlike prior methods that primarily focus on images and states as conditions, we pioneer in incorporating 3D conditioning into diffusion policies.
@@ II-B Visual Imitation Learning
Figure 2: Overview of 3D Diffusion Policy (DP3). Above: In the training phase, DP3 simultaneously trains its perception module and decision-making process in an end-to-end manner using expert demonstrations. During evaluation, DP3 determines actions based on visual observations from the environment. Below: DP3 perceives its environment through single-view point clouds. These are transformed into compact 3D representations by a lightweight MLP encoder. Subsequently, DP3 generates actions conditioning on these 3D representations and the robot’s states, using a diffusion-based backbone.
@@ II-C Learning Dexterous Skills
Achieving human-like manipulation skills in robots has been a longstanding objective pursued by robotics researchers. Reinforcement learning has been a key tool in this endeavor, enabling robots with dexterous hands to master a variety of tasks, such as pouring water [48, 82], opening doors [50, 21, 8], rotating objects [45, 77, 79, 46], reorienting objects [18, 7, 6], spinning pens [34], grasping tools [1], executing handovers [84, 24], and building Legos [9]. Imitation learning offers another pathway, with approaches like DIME [3] and DexMV [48] translating human hand movements into robotic actions through retargeting and enabling learning from human videos. Our work, however, diverges from these specific design-centric methods. We demonstrate that enabling the acquisition of these complex skills with minimal demonstrations could be achieved by improving the imitation learning algorithm itself.
@@ III-C Decision
Implementation details. We use the convolutional network-based diffusion policy [10]. We use DDIM [63] as the noise scheduler and use sample prediction instead of epsilon prediction for better high-dimensional action generation, with 100 timesteps at training and 10 timesteps at inference. We train 1000 epochs for MetaWorld tasks due to its simplicity and 3000 epochs for other simulated and real-world tasks, with batch size 128 for DP3 and all the baselines.
@@ IV-B Efficiency and Effectiveness
    High accuracy. Summarized results are in Figure 1(a) and results for each domain are in Table I. We observe that DP3 achieves a success rate exceeding 90% in nearly 30 tasks, whereas Diffusion Policy does in fewer than 15 tasks. Additionally, DP3 did not record any task with a success rate below 10%, in contrast to Diffusion Policy, which had more than 10 tasks below 10%. Note that most of the tasks are only trained with 10 demonstrations.
@@
    Learning efficiency. While we train all the algorithms for 3000 epochs to guarantee convergence, we observe that DP3 typically reaches convergence within approximately 500 epochs across all tasks, as illustrated in Figure 5. In contrast, Diffusion Policy tends to converge at a much slower pace or converge into sub-optimal results.
@@
    Efficient scaling with demonstrations. As shown in Figure 6, we find that in Adroit tasks, both DP3 and Diffusion Policy perform reasonably while DP3 achieves a comparable accuracy with fewer demonstrations. For some MetaWorld tasks above the easy level such as Assembly and Disassemble, DP3 could achieve higher accuracy when demonstrations are sufficient. This underscores that the 3D modality is not just beneficial but essential for certain manipulation tasks.
@@ IV-C Ablations
Choice of 3D representations. In DP3, we deliberately select point clouds to represent the 3D scene. To compare different choices of 3D representations, we implement other 3D representations, including RGB-D, depth, and voxel. We also compare with oracle states, which include object states, target goals, and robot velocity besides robot poses. The RGB-D and depth images are processed using the same image encoder as Diffusion Policy, while voxel representations employ the VoxelCNN, as implemented in [7]. As demonstrated in Table IV, these alternative 3D representations fall short of DP3. We note that RGB-D and depth images are close and not comparable to point clouds, indicating that the proper usage of depth information is essential. Additionally, we observe that point clouds and oracle states are very competitive, showing that point clouds might help learn an optimal policy from demonstrations.
@@
Oracle State $99\scriptstyle{\pm 2}$ $61\scriptstyle{\pm 2}$ $\mathbf{44\scriptstyle{\pm 3}}$ $94\scriptstyle{\pm 1}$ $\mathbf{72\scriptstyle{\pm 7}}$ $91\scriptstyle{\pm 8}$ $76.8$
Point cloud $\mathbf{100\scriptstyle{\pm 0}}$ $\mathbf{62\scriptstyle{\pm 4}}$ $43\scriptstyle{\pm 6}$ $\mathbf{99\scriptstyle{\pm 1}}$ $69\scriptstyle{\pm 4}$ $\mathbf{97\scriptstyle{\pm 4}}$ $\mathbf{78.3}$
@@
Design choices in DP3. Besides the 3D representations, the effectiveness of DP3 is contributed by several small design choices, as shown in Table VII. (a) Cropping point clouds helps largely improve accuracy; (b) Incorporating LayerNorm layers could help stabilize training across different tasks [22, 4]; (c) Sample prediction in the noise sampler brings faster convergence, also shown in Figure 7; (d) The projection head in DP3 Encoder accelerates the inference by projecting features to the lower dimension, without hurting accuracy; (e) Removing color channels ensures robust appearance generalization; (f) In low-dimensional control tasks, DPM-solver++ [33] as the noise sampler is competitive to DDIM, while DPM-solver++ could not handle high-dimensional control tasks well.
@@ V-C Generalization
View generalization. Generalizing image-based methods across different views is notably challenging [76], and acquiring training data from multiple views can be time-consuming and costly [83, 59]. We demonstrate in Table XII that DP3 effectively addresses this generalization problem when the camera views are altered slightly. It is important to note that since the camera view is altered, we manually transform the point clouds and adjust the cropped space of the point clouds. Accurate transformation isn’t necessary due to the robustness of our network. However, it is crucial to acknowledge that while the network can generalize across minor variations in camera views, significant changes might be hard to handle.
@@
Cluttered Scenes. Despite the simplicity of the DP3 Encoder, we demonstrate that DP3 is capable of handling tasks in complex real-world cluttered environments. To illustrate this, we design a pick & place task (i.e. pick the cube and place it in the bowl) set in cluttered scenes using a gripper and collect 50 demonstrations for training. The results presented in Table XIII show that DP3 solves the task with a high success rate. Aligning with our simulation experiments, DP3 equipped with PointNeXt fails to address the task. Meanwhile, DP3 using color point clouds as input performs comparably to the original DP3 when picking the training yellow cube, yet it struggles with other colored cubes and diverse objects. This demonstrates the effectiveness and generalization ability of DP3 in complex scenes.
TABLE XIII: Results in cluttered scenes. Each algorithm is evaluated with 10 trials in the training color. Each out-of-domain color and object are evaluated with one trial.
Cluttered Scenes Diffusion Policy DP3 w/ PointNeXt DP3 w/ color DP3
Success Rate $60$ 00 $\mathbf{80}$ $\mathbf{80}$
Train with $\blacksquare$ Cube $\blacksquare$ $\blacksquare$ $\blacksquare$ Charger Cylinder Rope
DP3 w/ color
DP3
V-D Observation on Deployment Safety
In our real-world experiments, we observe that image-based and depth-based diffusion policies often deliver unpredictable behaviors in real-world experiments, which necessitates human termination to ensure robot safety. We define this situation as safety violation and compute the safety violation rate in our main real-world experiments, shown in Table XIV. Interestingly and surprisingly, we find that DP3 rarely violates the safety, showing that DP3 is a practical and hardware-friendly method for real robot learning. An intuitive explanation is that since the robots operate in 3D space, directly observing 3D information helps avoid collision. It is important to note that our assessment of safety is primarily qualitative. We intend to explore a more theoretical understanding of this observation in our future work.
TABLE XIV: Safety violation rate. While conducting the main real-world experiments, we also count the times of safety violation and compute the rate.
@@ Acknowledgement
We would like to thank Zhecheng Yuan, Chen Wang, Cheng Lu, and Jianfei Chen for their helpful discussions. This work is supported by National Key R&D Program of China (2022ZD0161700).